# LAG: Logic-Augmented Generation from a Cartesian Perspective

### v2

@@ Introduction
Llama3 (8B) (Meta 2024) 23.7 20.1 33.8 15.4 6.4 6.0
Llama3 (13B) (Meta 2024) 17.7 11.1 22.3 4.7 2.3 2.0
GPT-3.5-turbo (OpenAI 2024) 31.5 35.4 31.0 29.9 7.9 10.9
GPT-4o-mini (OpenAI 2024) 30.4 34.2 29.0 28.6 7.8 10.1
@@
CoT (Top-5) (Wei et al. 2022) 55.1 57.1 48.7 45.9 27.1 30.7
IRCoT (Top-5) (Trivedi et al. 2023) 58.4 59.6 53.0 36.8 22.6 26.1
Novel Retrieval-Augmented-Generation Methods
KGP (Wang et al. 2024) 56.2 57.1 52.2 33.9 30.5 27.3
G-retriever  (He et al. 2024) 41.3 40.9 47.8 25.7 14.1 15.6
RAPTOR  (Sarthi et al. 2024) 58.1 55.3 60.6 43.9 32.2 29.7
LightRAG  (Guo et al. 2025) 61.5 60.5 54.4 38.0 27.7 28.3
HippoRAG (single-step) (Gutiérrez et al. 2024) 55.2 57.9 63.7 57.5 31.4 30.1
HippoRAG (multi-step) (Gutiérrez et al. 2024) 61.1 63.6 66.4 62.4 34.0 31.8
GFM-RAG (single-step)  (Luo et al. 2025) 61.4 64.8 66.2 61.1 29.3 32.6
GFM-RAG (multi-step) (Luo et al. 2025) 63.4 65.5 69.5 63.2 31.5 35.5
HippoRAG 2 (Gutiérrez et al. 2025) 61.2 64.3 62.0 58.8 34.5 35.6
LAG(Ours) 68.3 69.4 71.3 64.0 42.8 43.5
@@
Experiment
Experimental Setup
Dataset
@@
To evaluate the effectiveness of LAG, we conducted experiments on three standard public multi-hop QA benchmarks: HotpotQA (Yang et al. 2018), MuSiQue (Trivedi et al. 2022), and 2WikiMultiHopQA (2Wiki) (Ho et al. 2020). Following the evaluation protocol of HippoRAG, we used the same corpus for retrieval and sampled 1,000 questions(same with HippoRAG) from each validation set as our test queries. This setup ensures a fair, apples‑to‑apples comparison across methods. To further assess LAG’s reasoning capabilities, we also evaluated it on the recently released GraphRAG‑Bench (Xiao et al. 2025), which proved that LAG not only can generate correct answers but also maintains the rationality of the reasoning. In the Appendix C, we provide additional experimental results regarding efficiency and comprehensiveness.
Baseline
@@
We compared our approach against a diverse set of established baselines, grouped into three categories: 1) LLM-Only: Direct question answering using a large language model without any external retrieval. 2) Vanilla Retrieval-Augmented Generation: Integration of semantic retrieval with chain‑of‑thought prompting to guide the LLM’s generation. 3) State‑of‑the‑Art RAG Systems: Recent, high-performing retrieval‑augmented generation methods, including HippoRAG 2 (Gutiérrez et al. 2025), GFM‑RAG (Luo et al. 2025), HippoRAG (Gutiérrez et al. 2024), LightRAG (Guo et al. 2025), KGP (Wang et al. 2024), G‑Retriever (He et al. 2024), and RAPTOR  (Sarthi et al. 2024). A detailed description of each baseline is provided in the Appendix B.
Metrics
@@
Exact string matching can be overly stringent for multi-hop QA, as variations in casing, grammar, tense, or paraphrasing may cause a correct response to be marked wrong. Following existing work (Wang et al. 2025), we adopt two complementary metrics: 1) Contain-Match Accuracy: Measures whether the predicted answer contains the gold answer as a sub-string. This metric accommodates minor surface-form differences while still enforcing semantic correctness. 2) GPT-Evaluation Accuracy: An LLM‑based evaluation in which the model receives the question, the gold answer, and the prediction, then judges whether the prediction is semantically equivalent to the gold answer. These metrics together provide a balanced assessment of both surface‑level fidelity and deeper semantic correctness. For challenging reasoning task, we follow metric setting of the benchmark (Xiao et al. 2025).
Implementation Details
@@
Both our proposed method and all baselines utilize GPT-4o-mini as the default LLM. All experiments were executed on the RTX 4090 D. For all the methods, we use all-MiniLM-L6-v2 as the embedding model. For top-k parameters across methods, we set k = 5.
Main Results
As shown in Table 1, the base LLM exhibited weak performance when directly addressing these complex questions. When simple semantic retrieval and CoT prompting were incorporated, response quality improved notably. Performance across novel RAG methods exhibits variability: KGP, RAPTOR and LightRAG demonstrated improvements in some scenarios, but they did not consistently outperform the Vanilla RAG. In contrast, HippoRAG, GFM-RAG, and HippoRAG 2 consistently achieved significant performance gains across all three datasets.
Our proposed LAG method significantly outperformed all baseline approaches on both Contain-Acc and GPT-Acc metrics. Specifically, compared to the default LLM (GPT-4o-mini), LAG achieved absolute improvements of approximately 40 points in both Contain-Acc and GPT-Acc on the HotpotQA and 2Wiki datasets, with similar gains of around 30 points observed on the Musique dataset. Relative to existing RAG baselines, LAG maintained a significant advantage over most methods; even compared to the strong performers HippoRAG 2 and GFM-RAG, LAG’s superiority remained pronounced, particularly when handling the challenging Musique dataset. Overall, these results confirm that LAG not only elevates answer accuracy across diverse domains but also ensures stable performance where other RAG approaches struggle. The marked improvements highlight LAG’s superior logical capabilities in RAG.
@@
KGP (Wang et al. 2024) 58.7 42.2
G-retriever (He et al. 2024) 60.2 43.7
LightRAG (Guo et al. 2025) 60.5 43.8
GFM-RAG (Luo et al. 2025) 60.4 44.3
HippoRAG (Gutiérrez et al. 2024) 60.9 44.6
HippoRAG 2 (Gutiérrez et al. 2025) 59.8 43.7
RAPTOR (Sarthi et al. 2024) 60.8 45.5
LAG(Ours) 65.2 46.4
Table 2: Reasoning performance comparison among state-of-the-art baselines and LAG on GraphRAG-Bench. R score is used to evaluate the consistency between the generated rationales and the gold rationales. AR Score is an evaluation of generated answers based on the R score.
Challenging Reasoning Task
@@
Verification of the Importance of Logic
We hypothesize that the effectiveness of our proposed LAG framework derives fundamentally from the preservation of logic. To substantiate this claim, we first invoke the Cartesian principle, which establishes a theoretical foundation for the role of logic in reasoning systems. We then perform an empirical validation: in the case of a vanilla RAG, we observe that maintaining a logical order yields markedly better performance on complex tasks. To isolate the contribution of logical order within LAG, we introduce two controlled perturbations: (1) We don’t concatenate the embedding of the former dependency question to the subsequent question. (2) We shuffle the order of the logical chain before retrieval and generation. Both interventions incur a statistically significant drop in performance, thereby confirming that logic is essential to RAG systems.
LAG HotpotQA 2Wiki Musique
GPT-4o-mini (Default LLM) 30.4 34.2 29.0 28.6 7.8 10.1
GPT-4o-mini + retrieval 54.6 56.8 46.6 45.3 25.6 29.0
w/ decompose 60.0 59.9 57.5 46.9 30.0 30.8
w/ decompose+ reorder 62.3 61.1 59.4 48.6 33.8 32.1
w/ decompose+ reorder+ chain reasoning 65.8 64.8 66.9 60.2 36.6 38.9
w/ decompose+ reorder+ chain reasoning+ terminator 68.3 69.4 71.3 64.0 42.8 43.5
Table 4: Ablation study on three standard datasets.
Question When did the explorer reach the city where the headquarters of the only group larger than Vilaiyaadu Mankatha’s record label is located?
Ground Truth August 3, 1769
Support Context [“Vilaiyaadu Mankatha”, “Sony Music”, “The Right Stuff Records”, “Santa Monica”]
Vanilla RAG Directly retrieved context:
1) ✕ “Francisco de Orellana”: Francisco de Orellana was a Spanish explorer…
2) ✕ “Malta Summit”: The meetings took place in the Mediterranean, off the…
3) ✓ “Vilaiyaadu Mankatha”: …and were marketed by Sony Music Entertainment…
4) ✕ “Military history of Indonesia”: Many inscriptions were found that mention …
5) ✕ “Menteng, Menteng”: Menteng is an administrative village in the Menteng …
Prediction:
✕ Guayaquil.
LAG Logical chain:
[…]
Sub-Q2: What company is the only group larger than #1? $\to$ Universal Music Group
Sub-Q3: Where is the headquarters of #2? $\to$ Santa Monica
Sub-Q4: What date did the explorer reach #3? $\to$ August 3, 1769
Retrieved context by logical chain:
1) ✓ “Vilaiyaadu Mankatha”: …and were marketed by Sony Music Entertainment
2) ✓ “Sony Music”: …second largest recorded music company, after Universal Music Group….
3) ✓ “The Right Stuff Records”: …Universal Music Group and is based out of Santa Monica
4) ✓ “Santa Monica”: …who camped near the present day intersection of … on August 3, 1769
Prediction:
✓ August 3, 1769.
Table 5: Case study of LAG. For complex questions, LAG yields solutions more consistent with human scientific reasoning.
Ablation Study
@@
To validate the effectiveness of each component in our proposed LAG, we conducted a ablation study. Results are presented in Table  4, with key observations as follows: The LLM-only baseline exhibited suboptimal performance in complex QA tasks. However, incorporating the retrieval module produced significant improvements, demonstrating the critical role of external knowledge retrieval. Specifically, adding the decomposition module further enhanced performance; we attribute to its ability to break down complex questions into simpler sub-questions, which facilitates more targeted and effective retrieval. Furthermore, integrating the reordering module led to additional gains by strengthening the logical coherence among sub-questions, optimizing the reasoning sequence. A more substantial performance boost was observed when introducing the core “logical chain reasoning” module, particularly in high-difficulty scenarios. This highlights the indispensable role of structured logical chains in guiding complex QA processes. Notably, incorporating the logical terminator module achieved the best overall performance. This improvement stems from its ability to mitigate error propagation in chain-based reasoning by terminating erroneous inference paths in a timely manner, thereby preventing cumulative errors.
Case Study
To demonstrate the superiority of LAG more intuitively, we present case study results in Table  5. For real-world complex questions, vanilla RAG exhibit critical limitations: simply performing retrieval based solely on the semantic information of the original question fails to retrieve useful knowledge. Additionally, LLMs may struggle to fully comprehend the question itself, resulting in a substantial discrepancy between predicted and ground-truth answers. In contrast, LAG addresses these challenges through a structured reasoning process: 1) decomposing the complex question into atomic sub-questions; 2) reordering these sub-questions according to logical reasoning sequences; 3) performing reasoning on sub-questions in this logical order, where retrieval and inference for each sub-question leverage answer information from previously solved, logically dependent sub-questions. This approach not only enables effective retrieval of relevant information but also supports reasonable inference for each step. This process adheres to a scientific reasoning paradigm consistent with Cartesian principles. It involves breaking down complex questions into manageable components and reasoning sequentially based on prior results, ultimately leading to the correct answer.
Conclusion

### v3

@@ 4.3 Integrated Generation
Llama3 (8B) Meta (2024) 23.7 20.1 33.8 15.4 6.4 6.0
GPT-3.5-turbo OpenAI (2024) 31.5 35.4 31.0 29.9 7.9 10.9
GPT-4o-mini OpenAI (2024) 30.4 34.2 29.0 28.6 7.8 10.1
@@
CoT (Top-5) Wei et al. (2022) 55.1 57.1 48.7 45.9 27.1 30.7
IRCoT (Top-5) Trivedi et al. (2023) 58.4 59.6 53.0 36.8 22.6 26.1
Novel Graph Retrieval-Augmented-Generation
KGP Wang et al. (2024) 56.2 57.1 52.2 33.9 30.5 27.3
G-retriever  He et al. (2024) 41.3 40.9 47.8 25.7 14.1 15.6
RAPTOR  Sarthi et al. (2024) 58.1 55.3 60.6 43.9 32.2 29.7
LightRAG  Guo et al. (2025) 61.5 60.5 54.4 38.0 27.7 28.3
HippoRAG (single-step) Gutiérrez et al. (2024) 55.2 57.9 63.7 57.5 31.4 30.1
HippoRAG (multi-step) Gutiérrez et al. (2024) 61.1 63.6 66.4 62.4 34.0 31.8
GFM-RAG (single-step)  Luo et al. (2025) 61.4 64.8 66.2 61.1 29.3 32.6
GFM-RAG (multi-step) Luo et al. (2025) 63.4 65.5 69.5 63.2 31.5 35.5
HippoRAG 2 Gutiérrez et al. (2025) 61.2 64.3 62.0 58.8 34.5 35.6
LinearRAG  Zhuang et al. (2025) 64.2 64.9 69.5 62.6 32.8 36.9
[…]
@@
5 Experiment
5.1 Experimental Setup
Dataset.
@@ 5.1 Experimental Setup
To evaluate the effectiveness of LAG, we conducted experiments on three standard public multi-hop QA benchmarks: HotpotQA Yang et al. (2018), MuSiQue Trivedi et al. (2022), and 2WikiMultiHopQA (2Wiki) Ho et al. (2020). Following the evaluation protocol of HippoRAG, we used the same corpus for retrieval and sampled 1,000 questions(same with HippoRAG) from each validation set as our test queries. This setup ensures a fair, apples‑to‑apples comparison across methods. To further assess LAG’s reasoning capabilities, we also evaluated it on the recently released GraphRAG‑Bench Xiao et al. (2025), which proves that LAG not only generate correct answers but also maintains the rationality of reasoning.
Baseline.
@@
We compared our approach against a diverse set of established baselines, grouped into three categories: 1) LLM-Only: Direct question answering using a large language model without any external retrieval. 2) Vanilla RAG: Integration of semantic retrieval with chain‑of‑thought prompting to guide the LLM’s generation. 3) SOTA GraphRAG Methods: Recent, high-performing GraphRAG methods, including HippoRAG 2 Gutiérrez et al. (2025), GFM‑RAG Luo et al. (2025), LinearRAG Zhuang et al. (2025), HippoRAG Gutiérrez et al. (2024), LightRAG Guo et al. (2025), KGP Wang et al. (2024), G‑Retriever He et al. (2024), and RAPTOR  Sarthi et al. (2024). Detailed description of baselines is in the Appendix.
Metrics.
@@
Exact string matching can be overly stringent for multi-hop QA, as variations in casing, grammar, tense, or paraphrasing may cause a correct response to be marked wrong. Following existing work Zhuang et al. (2025), we adopt two complementary metrics: 1) Contain-Match Accuracy: Measures whether the predicted answer contains the gold answer as a sub-string. This metric accommodates minor surface-form differences while still enforcing semantic correctness. 2) GPT-Evaluation Accuracy: An LLM‑based evaluation in which the model receives the question, the gold answer, and the prediction, then judges whether the prediction is semantically equivalent to the gold answer. These metrics together provide a balanced assessment of both surface‑level fidelity and deeper semantic correctness. For challenging reasoning task, we follow metric setting of the benchmark Xiao et al. (2025).
Implementation Details.
@@
Both our proposed method and all baselines utilize GPT-4o-mini as the default LLM. All experiments were executed on the RTX 4090 D. For all the methods, we use all-MiniLM-L6-v2 as the embedding model. For top-k parameters across methods, we set k = 5. In the Appendix, we provide additional experimental results regarding efficiency, hyperparameters, LLM backbones and comprehensiveness.
@@
KGP Wang et al. (2024) 58.7 42.2
G-retriever He et al. (2024) 60.2 43.7
LightRAG Guo et al. (2025) 60.5 43.8
GFM-RAG Luo et al. (2025) 60.4 44.3
HippoRAG Gutiérrez et al. (2024) 60.9 44.6
HippoRAG 2 Gutiérrez et al. (2025) 59.8 43.7
RAPTOR Sarthi et al. (2024) 60.8 45.5
LinearRAG Zhuang et al. (2025) 61.5 45.4
[…]
Table 2: Reasoning performance comparison among SOTA baselines and LAG on GraphRAG-Bench. R score is used to evaluate the consistency between the generated rationales and gold rationales. AR Score is an evaluation of generated answers based on the R score.
Figure 3: Ablation study of proposed LAG on three standard datasets.
5.2 Main Results
As shown in Table 1, the base LLM exhibited weak performance when directly addressing these complex questions. When simple semantic retrieval and CoT prompting were incorporated, response quality improved notably. Performance across novel RAG methods exhibits variability: KGP, RAPTOR and LightRAG demonstrated improvements in some scenarios, but they did not consistently outperform the Vanilla RAG. In contrast, HippoRAG, LinearRAG, GFM-RAG, and HippoRAG 2 consistently achieved performance gains across all datasets.
Our proposed LAG method significantly outperformed all baseline approaches on both Contain-Acc and GPT-Acc metrics. Specifically, compared to the default LLM (GPT-4o-mini), LAG achieved absolute improvements of approximately 40 points in both Contain-Acc and GPT-Acc on the HotpotQA and 2Wiki datasets, with similar gains of around 30 points observed on the MuSiQue dataset. Relative to existing RAG baselines, LAG maintained a significant advantage over most methods; even compared to the strong performers LinearRAG, HippoRAG 2 and GFM-RAG, LAG’s superiority remained pronounced, particularly when handling the challenging MuSiQue dataset. Overall, these results confirm that LAG not only elevates answer accuracy across diverse domains but also ensures stable performance where other RAG approaches struggle. The marked improvements highlight LAG’s superior logical capabilities in RAG.
5.3 Challenging Reasoning Task
@@
5.4 Verification of the Importance of Logic
We hypothesize that the effectiveness of our proposed LAG derives fundamentally from the preservation of logic. To substantiate this claim, we first invoke the Cartesian principle, which establishes a theoretical foundation for the role of logic in reasoning systems. We then perform an empirical validation: in the case of a vanilla RAG, we observe that maintaining a logical order yields markedly better performance on complex tasks. To isolate the contribution of logical order within LAG, we introduce two controlled perturbations: (1) We don’t concatenate the embedding of the former dependency question to the subsequent question. (2) We shuffle the order of the logical chain before retrieval and generation. Both interventions incur a statistically significant drop in performance, thereby confirming that logic is essential to RAG systems.
@@ 5.4 Verification of the Importance of Logic
5.5 Ablation Study
@@
To validate the effectiveness of each component in our proposed LAG, we conducted a ablation study. Results are presented in Figure  3, with key observations as follows: The LLM-only baseline exhibited suboptimal performance in complex QA tasks. However, incorporating the retrieval module produced significant improvements, demonstrating the critical role of external knowledge retrieval. Specifically, adding the decomposition module further enhanced performance; we attribute to its ability to break down complex questions into simpler sub-questions, which facilitates more targeted and effective retrieval. Furthermore, integrating the reordering module led to additional gains by strengthening the logical coherence among sub-questions, optimizing the reasoning sequence. A more substantial performance boost was observed when introducing the core “logical chain reasoning” module, particularly in high-difficulty scenarios. This highlights the indispensable role of structured logical chains in guiding complex QA processes. Incorporating the atomic memory bank can further enhance the accuracy. Notably, incorporating the logical terminator module achieved the best overall performance. This improvement stems from its ability to mitigate error propagation in chain reasoning by terminating erroneous inference paths in a timely manner, thereby preventing cumulative errors. These findings confirm that each module adds unique value, necessitating their integration.
6 Conclusion
@@
Limitations
While LAG has advanced the paradigm for text-based retrieval-augmented generation, it currently lacks support for multimodal input sources. Extending this framework to incorporate multimodal inputs would enable a more comprehensive modeling of human information processing and reasoning mechanisms involving multimodal data. Given that real-world information is inherently multimodal, such an extension would further enhance the applicability and robustness of the LAG model.
Ethics Statement
We confirm that we have strictly adhered to the ACL Ethics Policy throughout this study. Our research employs four publicly available datasets: HotpotQA, 2WikiMultiHopQA, MuSiQue, and GraphRAG-Bench. HotpotQA, 2WikiMultiHopQA, and MuSiQue are developed to assess the complex question-answering capabilities of various models, while GraphRAG-Bench is a domain-specific reasoning benchmark tailored for retrieval-augmented generation methods. All datasets utilized in this work have been widely adopted in RAG-related research and are free of private, sensitive, or personally identifiable information. We carefully selected these datasets to ensure compliance with ethical standards and to mitigate potential biases. Notably, our study does not involve the collection or modification of user-generated content, nor do we introduce synthetic data that may give rise to unintended misinformation.